CAN ONE DETERMINE THE ELECTRIC DIPOLAR MOMENT OF ArH^{+} FROM FREQUENCY DATA? \underline{J. F. OGILVIE}, Centre for Experimental and Constructive Mathematics, Department of Mathematics, Simon Fraser University, 8888 University Drive, Burnaby, British Columbia V5A 1S6 Canada.